Opened 4 years ago

Closed 3 years ago

#591 closed flight processing (fixed)

RG13/08, flight day 234b/2015, Monks Wood

Reported by: lah Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2015/flight_data/uk/RG13_08-2015_234b_Monks_Wood

Data arrived from ARSF via USB on 27/08/2015.

Scientific objective: Measuring ash die back.

Priority: Unknown

PI: David Coomes

Flown on same day as Cambridge (RG13/08) and Alconbury(GB15/00).

Sensors:

  • Fenix
  • Leica LIDAR
  • FW Leica LIDAR
  • RCD

Change History (33)

comment:1 Changed 4 years ago by lah

Unpacking

All extra files have been split out except from RCD, which is duplicated for 234a, 234b and 234c. Please only process RCD for one of these projects, then split once processed.

comment:2 Changed 4 years ago by gej

RCD processing
Have produced tif files, waiting to do navigation.

comment:3 Changed 4 years ago by stgo

Nav processing

lat 52 11 07.32948
long -0 06 44.97121
height 120.043
Version 0, edited 4 years ago by stgo (next)

comment:4 Changed 4 years ago by stgo

Nav processing

Navigation data processed with seperation below 5 cm variation and a good position accuracy. Proceeding to lidar processing.

comment:5 Changed 4 years ago by stgo

Lidar processing

There were only 3 lines and all are seperated so much overlap info cannot be estimated and pitch/roll/heading cant be calculated. Have generated a delivery from the data (there is the same amount of data for hyperspectral so I assume its safe to proceed with just three lines)

comment:6 Changed 4 years ago by stgo

Lidar processing

Completed delivery. ready for delivery check

comment:7 Changed 4 years ago by lah

Fenix Processing

Found SCTs:

Flightline FENIX
7 1.00
8 -0.02
9 1.00

comment:8 Changed 4 years ago by lah

Fenix Processing
Delivery created, ready for checking.

comment:9 Changed 4 years ago by lah

Lidar Delivery
Replaced logsheet with fixed version.

comment:10 Changed 4 years ago by lah

Lidar DC

  • Not many buildings, but lines up very well with 2014 152a
  • Slightly too many returns classified as noise (generally second returns in forested areas). Could be important for application?
  • changed las1.0 to 1.2
  • changed sol file name -234b to _234b
  • demcompare mean -0.94
  • fw look as normal

Ready to go if lots of noisy points not a problem.

comment:11 Changed 4 years ago by gej

RCD processing

Have created camera navigation file, tagging tiff images now.

comment:12 Changed 4 years ago by gej

RCD processing

Delivery and Readme created, ready for delivery check

comment:13 Changed 4 years ago by stgo

Fenix delivery check

starting check

Last edited 4 years ago by stgo (previous) (diff)

comment:14 follow-up: Changed 4 years ago by stgo

Fenix delivery check

A couple of things.

  • apl command check fails on mapping stage. This is because the igm filename is incorrect (3b vs 1b), I'm checking the generator and checking script to see if this is a problem with them.
  • should we be worried about proj tidy saying "The number of wavelengths in FENIX234-15-8.hdr is not the same as the length as any of the .wls files in /users/rsg/arsf/calibration/2015/fenix/" ? (says this for all files)

comment:15 in reply to: ↑ 14 ; follow-up: Changed 4 years ago by lah

Replying to stgo:

Fenix delivery check

A couple of things.

  • apl command check fails on mapping stage. This is because the igm filename is incorrect (3b vs 1b), I'm checking the generator and checking script to see if this is a problem with them.

There's not a problem with the names generated by the script; these were added manually. The problem is that if there is no line 1 it doesn't generate the commands automatically so human error is far more likely.

  • should we be worried about proj tidy saying "The number of wavelengths in FENIX234-15-8.hdr is not the same as the length as any of the .wls files in /users/rsg/arsf/calibration/2015/fenix/" ? (says this for all files)

That is odd, as the wavelengths are from 201509. Are you sure this isn't just for the original_headers subdirectory?

comment:16 in reply to: ↑ 15 Changed 4 years ago by stgo

Replying to lah:

That is odd, as the wavelengths are from 201509. Are you sure this isn't just for the original_headers subdirectory?

I've just checked, it only looks at the files in the sensor directory and won't go down into the original headers folder. APL would have complained before this point so I *think* it's ok.

I'd assumed the apl commands were auto generated, they can be fixed by changing igm file references from 3b to 1b.

comment:17 Changed 4 years ago by dac

RCD Delivery Check

Starting delivery check.

comment:18 Changed 4 years ago by dac

RCD Delivery Check

All looks fine - ready to deliver.

comment:19 Changed 4 years ago by lah

Fenix Delivery

Have corrected filenames in config file and regenerated Readme. Should all be ok now.

comment:20 Changed 4 years ago by stgo

Lidar delivery

Recreated the delivery with much reduced noisey points, would be worth a quick check of readme but otherwise are identical files

comment:21 Changed 4 years ago by lah

Lidar re-DC

  • classification looks much better
  • renamed las1.0 to 1.2
  • changed sol file name in delivery again

Everything else looks fine, so ready to deliver.

comment:22 Changed 4 years ago by stgo

Fenix delivery

Can't see anything wrong with the delivery now. Need someone to double check spectra but otherwise ready for delivery.

comment:23 Changed 4 years ago by dac

Fenix Delivery Check

Checked spectra against Py6S. Looks fine - ready to delivery.

comment:24 Changed 4 years ago by dac

Fenix

Started zipping mapped files ready for delivery.

comment:25 Changed 4 years ago by dac

Fenix, LiDAR and RCD Delivery

Sent to PI via FTP (arsf19) 17/12/2015

comment:26 Changed 4 years ago by stgo

LiDAR processing

With the updated processing reg the overlap has been significantly increased, as was stated before the data lines up with 152 and what features are available to estimate from do not seem to show any alterations in pitch and roll.

Creating delivery.

comment:27 Changed 4 years ago by stgo

LiDAR processing

Created lidar delivery, ready for check.

comment:28 Changed 4 years ago by dac

LiDAR Delivery Check

Starting delivery check

comment:29 Changed 4 years ago by dac

LiDAR Delivery Check

Looks fine. The .trj files are missing, not sure if these need to be generated again or ones from previous delivery can be used.

The quality of the ASTER DSM is really poor - it is very noisy. I was unsure if there was a problem with the LiDAR so also compared to SRTM and there was a good match there so I think it is just a problem with the ASTER.

Once .trj files are added ready to deliver.

comment:30 Changed 4 years ago by dac

LiDAR

Re-generated .trj files. The .reg file was missing so needed to re-create - will check with stgo (likely local copy).

Ready to deliver

comment:31 Changed 4 years ago by dac

LiDAR data delivery

Redelivered via FTP (arsf19) - 26/02/2016

comment:32 Changed 4 years ago by dac

Archiving

Started uploading to NEODC.

comment:33 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC. Closing ticket.

Note: See TracTickets for help on using tickets.